A 65.0mL aliquot of a 1.30M solution is diluted to a total volume of 208 mL. A 104 mL portion of that solution is diluted by adding 173 mL of water. What is the final concentration? Assume the volumes are additive.
Answer #1

By dilution law,

M1V1 = M2V2

1.30×65.0 = 208×M2

M2 = 0.406 M

In 2nd case

M2V3 = M3V4

0.406 × 104 = M3×(173+104)

M3 = 0.153 M